Air is escaping from a spherical balloon at the rate of 2π‘π‘š3 per minute. How fast is the radius shrinking when the volume is 36πœ‹ π‘π‘š3 ?


Find the rate of change of the area 𝐴, of the circle with respect to its circumference C, 𝑖. 𝑒 𝑑𝐴 𝑑�


1
Expert's answer
2021-04-25T17:41:20-0400

"Volume={4 \over 3} \pi r^3"

Since volume is 36πœ‹ π‘π‘š3

Then, "36\pi= {4 \over 3}\pi r^3"

"r^3=27"

r=3cm


Since, "V={4 \over 3} \pi r^3" and "{dv \over dt}=" 2cm3 per mins

"{dv \over dt}=4\pi r^2{dr \over dt}"


"2=4\pi(3^2) {dr \over dt}"


"{dr \over dt}={1 \over 18\pi}cm/mins"




"Area(A)=\pi r^2"


"Circumference (C)=

2\pi r"

"r={C \over 2\pi}"


"\therefore 
A= \pi({C \over 2\pi})^2"


"A= {C^2 \over 4\pi}"


"{dA \over dC}={C \over 2\pi}units"